HUBBLE SPACE TELESCOPE DAILY REPORT #1886 PERIOD COVERED: 0000Z (UTC) 05/27/97 - 0000Z (UTC) 05/28/97 Daily Status Report as of 148/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Five Sets of NIC /2/3 7048 (NICMOS Parallel Thermal Background Test) The Near Infrared Cameras (Cameras Two and Three) were used to make observations of an object designated as Target-1 in order execute measurements of the stability of the HST thermal emissions. The observations were executed as scheduled, and no problems were reported. 1.2 Completed Five Sets of WF/PC-2 6909 (WF/PC-2 Cycle 6 Calibration: Earth Flats) The WF/PC-2 was used to make observations of the bright Earth to obtain streak flats which are used to construct high quality flat fields for the WF/PC-2 filter set. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed Two Sets of STIS/CCD 7095 (STIS CCD Vignetting, Spatial Resolution in Spectroscopic Modes) The Space Telescope Imaging Spectrograph (CCD) was used to make observations of the point source P041C in order to study the vignetting and special resolution of the CCDs in the spectroscopic mode. This is to measure STIS relative response to a point source in the first order modes along the slit direction. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed WF/PC-2 6904 (WF/PC-2 Cycle 6 Calibration: Darks) The WF/PC-2 was used to make a series of several dark calibration images as part of a standard calibration within Cycle 6. This calibration procedure is performed once per week as part of an on going CCD calibration to be used to monitor and characterize hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed Four Sets of WF/PC-2 7553 (Pure Parallel Archive Survey II) WF/PC-2 parallel observations were obtained in the ultraviolet in search for interesting objects. This program is designed to find extragalactic sources with near ultraviolet excesses (3000 A0) outside the galactic plane and far ultraviolet sources (1600 A0) in the galactic plane. The procedure was executed as scheduled, and no anomalies were reported. 1.6 Completed Three Sets of WF/PC-2 7203 (Cosmology with the Medium Deep Survey - Short Orbit Cases) The WF/PC-2 was used to take short (about 20 minute) exposures, to be taken in parallel with the prime science observations, using the I filter. These exposures are part of the medium-deep survey, which emphasizes exposures that will be useful in obtaining cosmological parameters. The observations were completed as planned, and no problems were reported. 1.7 Completed STIS/CCD 7092 (CCD Dark Rate and Read Noise) Observations were taken to measure CCD dark current and gather data for making super darks. The observations were completed as planned, and no anomalies were reported. 1.8 Completed STIS/MA2 7093 (Dark Measurement for MAMA Band 1, and Band 2) The Space Telescope Imaging Spectrograph (MA2) was used to make dark calibration observations of the MAMA Band 1 and Band 2. These observations are to be used to establish that the detectors are performing nominally, that no internal movement of the MCP has taken place, and to establish the dark level in orbit for a given operating temperature. The observations were completed as planned, and no anomalies were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 3 Successful : 3 Scheduled Reacquisitions : 10 Successful : 10 2.2 FHST Updates: Scheduled : 4 Successful : 4 2.3 Operations Notes: An SSR interrupt log dump was performed at 146/2311Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science and SMOV operations. /DMH